OpenStudy (order115):

Statistics P values --- probability P is observed to be 0,55. We're testing H0: P=0.5 against H1: P>0.5. What is the p-value for this test? Answer is 0.16: How to get there?

OpenStudy (order115):

@mww and @Gokuporter

OpenStudy (mww):

Your hypothesis test. Null hypothesis. Probability is 50/50 (0.50) Alternative hypothesis P > 0.5 You statistic for a proportion is then given by \[z = \frac{ p - p_0 }{ \sqrt{\frac{ p(1-p) }{ n }} }\] where p is the actual probability, p0 is the assumed probability according to the null hypothesis and n is the number of observations. Plug these in, then use a set of z-tables to compute the corresponding p-value greater than the z-score obtained.

OpenStudy (order115):

Thanks! I managed to get the answer in the end :)